OpenStudy (anonymous):

Prove ||u+v||^2 + ||u-v||^2=2||u||^2+2||v||^2 anybody got an idea where I can start? I can't seem to remember the property I can use to simplify ||u+v||^2

OpenStudy (anonymous):

start with ||u+v||^2 = <u+v> dot <u+v>

OpenStudy (anonymous):

<u+v> dot <u+v> = ??

OpenStudy (anonymous):

and then <u-v> dot <u-v> = ??? add the two and you'll get what you trying to show

OpenStudy (anonymous):

Oh so just expand the equation. Thank you!

OpenStudy (anonymous):

<u+v> dot <u+v> = ||u||^2 + ||v||^2 + 2 u dot v